In a survey of 79 people, 40 people said they liked eggs, 37 people said they liked toast, and 18 people said they liked neither. How many people like both eggs and toast?
A) 16 B) 61 C) 3 D) 48

Answer by ikleyn(52754)

.
From the condition, the number of those who like at least one of the two foods is 79 - 18 = 61.


We can calculate this number by another way

61 = 40 + 37 - x,   (1)


where x is the number of those who like both foods.


Why we subtract x from the sum 40 + 37 ?  - Due to that simple reason that we count twice those who like both the foods when we add 40 and 37.


Then from (1),  x = (40+37) - 61 = 16.


Answer.  Under the given condition, the number of people who like both eggs and toasts is 16.


For similar solved problems see the lesson
    - Counting elements in sub-sets of a given finite set
in this site.